# An asymptotic stability proof and a port-Hamiltonian physics-informed neural network approach to chaotic synchronization

### v2

@@ 4 Numerical simulations and discussion
To corroborate the analytical results of Theorem 3.2 and the energy-based analysis of Section 3.2, we performed numerical simulations of the coupled memristive HR neuron system in Eq. (2) and its corresponding synchronization-error dynamics in Eq. (6). The two neurons were initialized with slightly different values in all five state variables to induce synchronization transients. We then computed the time evolution of the error variables $(e_{x},e_{y},e_{z},e_{u},e_{\phi})$, together with the Lyapunov and Hamilton energy derivatives $\dot{V}$ and $\dot{H}$, and analyzed their dependence on the parameters $m$, $k$, $\rho$, and $g_{e}$.
Figure 5 shows the temporal evolution of the five synchronization–error components, each representing the difference between corresponding variables of the two coupled neurons. All error trajectories decay asymptotically to zero after a brief transient, confirming CS. The fast variables $e_{x},e_{y},e_{u},e_{\phi}$ converge earlier than the slow adaptation variable $e_{z}$, reflecting the intrinsic separation of time scales in the neuron dynamics. These results corroborate the theoretical prediction that the synchronization manifold $\mathcal{M}=\{x_{1}=x_{2},\ldots,\phi_{1}=\phi_{2}\}$ is globally asymptotically stable under the given parameter values and coupling conditions.
Figure 5: (a)-(e): Time evolution of the five synchronization-error components $(e_{x},e_{y},e_{z},e_{u},e_{\phi})$, respectively, showing convergence to zero.
Figure 6 presents the temporal evolution of the synchronization Hamiltonian $H(e_{x},e_{y},e_{z},e_{u},e_{\phi})$ and its derivative $\dot{H}$ associated to the errors in Fig. 5. The Hamiltonian decays monotonically to zero as the neurons approach synchrony. The derivative $\dot{H}$ exhibits transient oscillations, taking both negative and positive values before stabilizing near zero; this reflects intermittent energy exchange due to memristive and flux-related terms in the dynamics. After transients, $\dot{H}\to 0$ and the system settles on the synchronization manifold, consistent with an overall dissipative relaxation toward synchrony.
Figure 6: Time evolution of the error‐system Hamiltonian function (a) and its time derivative (b), corresponding to the errors shown in Figure 5; as the system synchronizes, both energy and its derivative approach zero.
@@
Figure 8: Heatmaps of the time derivatives of the Lyapunov function (a1)–(a3) and the corresponding Hamiltonian function (b1)–(b3) as functions of $g_{e}$ and the system parameters $k$, $m$, and $\rho$, respectively. Across all parameter pairs, $\dot{V}$ and $\dot{H}$ display consistent qualitative trends, demonstrating a robust correspondence between Lyapunov-based and Hamiltonian-based stability indicators.
In summary, the numerical results confirm the analytical predictions of Section 3. The coupled neurons exhibit global asymptotic CS, with the Hamiltonian energy decaying monotonically to zero and the Lyapunov and Hamiltonian formulations yielding consistent measures of stability and energetics. These findings provide a basis for the pH-PINN framework introduced in Section 5, which exploits the equivalence between Lyapunov and Hamiltonian descriptions to learn the underlying energy structure directly from data, bridging analytical theory with machine-learning discovery of energy-conserving and dissipative dynamics.
@@ 5 A port-Hamiltonian physics-informed neural network (pH-PINN) approach
To study synchronization energy of the coupled HR neurons in Eq. (2) from a data-driven perspective, we employ a pH–PINN. This framework is well-suited to our system, whose synchronization behavior admits a Hamiltonian energy structure consistent with the pH formulation of conservative and dissipative flows. Analytical Hamiltonian approaches can establish stability but often become infeasible for high-dimensional or partially known systems. In such settings, the governing physical laws, such as energy conservation and dissipation, remain valid even when the explicit equations are unknown. The pH–PINN incorporates these physical constraints into the learning process, allowing the synchronization energy and its dynamics to be inferred directly from data. Here, the known analytical energy serves as a benchmark to quantitatively validate the learned representation. In the sequel, we briefly outline the principles of PINNs, HNNs, and pHNNs, and describe how these are integrated to construct the proposed pH-PINN framework.
@@ 5.2 Hamiltonian neural networks (HNNs)
HNNs [35] employ a scalar-valued network $H_{\theta}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\!\to\!\mathbb{R}$ that maps the state vector $\mathbf{x}$ to an energy $H_{\theta}(\mathbf{x})$. Automatic differentiation provides the gradient $\nabla H_{\theta}(\mathbf{x})$, representing partial derivatives of $H_{\theta}$ with respect to the state components. Rather than predicting the vector field directly, the dynamics are recovered through a (possibly state-dependent) skew-symmetric matrix $\mathbf{J}(\mathbf{x})\in\mathbb{R}^{n\times n}$:
@@ 5.3 Port-Hamiltonian neural networks (pHNN)
Here, $\mathbf{J}$ encodes power-conserving interconnections, and $\mathbf{R}\in\mathbb{R}^{n\times n}$ represents a non-conservative part of the dynamics. The energy rate satisfies
@@ 5.4 Port–Hamiltonian physics-informed neural network (pH-PINN)
Each term enforces a distinct physical law: $\mathcal{L}_{\mathrm{\tiny DynRes}}$ ensures dynamic consistency; $\mathcal{L}_{\tiny J}$ and $\mathcal{L}_{\tiny R}$ align learned and analytical conservative/non-conservative flows; $\mathcal{L}_{\mathrm{\tiny Cons}}$ enforces energy invariance along conservative trajectories; and $\mathcal{L}_{\tiny{\dot{H}}}$ enforces the energy–rate identity (Eq. 30).
The synchronization energy is parameterized by a multilayer perceptron (MLP) $f_{\theta}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\!\to\!\mathbb{R}$ and anchored at the synchronization manifold to remove gauge freedom and ensure a strict local minimum:
@@
This structure, inferred from $\mathbf{F}_{c}$ and $\mathbf{F}_{d}$ (Eq. (21)), reveals a sparse coupling pattern. We thus mask $\mathbf{J}_{\!\phi}$ and constrain $\mathbf{R}_{\!\psi}$ to be diagonal, consistent with standard pH learning [39]. Because the sparsity pattern depends on the Hamiltonian gradient $\nabla H$, we adopt a separable-gradient assumption, noting that many physical systems possess approximately separable energy functions [35, 57, 37]. Each component $\partial H/\partial e_{i}$ thus depends primarily on its own coordinate (and possibly a small subset), consistent with the weak cross-partials observed in $\mathbf{F}_{c}$ and $\mathbf{F}_{d}$. Under this assumption, rows 2255 of $\mathbf{F}_{c}$ depend only on $e_{x}$, while $F_{c,x}$ depends on $e_{y},e_{u},$ and $e_{\phi}$ (but not $e_{z}$), motivating a masked, skew-symmetric $\mathbf{J}_{\!\phi}$ ($\mathbf{J}_{\!\phi}^{\!\top}=-\mathbf{J}_{\!\phi}$). Inspection of $\mathbf{F}_{d}$ further reveals primarily self-coupling (coordinate-wise damping) terms, motivating a diagonal parameterization of $\mathbf{R}_{\!\psi}$.
@@ 5.5 Training and Results
We train the pH–PINN by minimizing the total loss in Eq. (31). Supervision uses raw derivatives from the linearized error system (Eq. (6)) and from the coupled HR system (Eq. (2)). Each sample comprises the error state $\mathbf{e}\in\mathbb{R}^{5}$, its ground–truth derivative $\dot{\mathbf{e}}^{\mathrm{true}}$, the full neuron states of both units (used only to build a short context vector $\mathbf{c}=(x_{1},u_{1},u_{2},\phi_{1})^{\top}\in\mathbb{R}^{4}$), and the analytical Hamiltonian and energy rate $(H,\dot{H})$ from Eqs. (24)–(25) for monitoring. We shuffle and split the data $80/20$ into train/validation; no normalization is used. All networks consume the concatenated $\mathbf{e},\mathbf{c}]$ as input. Implementation details are summarized in Table 1, and the overall architecture is shown in Fig. 9.
@@
Figure 9: Schematic diagram of the proposed pH-PINN architecture.
We train using the AdamW optimizer [58] with a linear learning-rate decay from $1\times 10^{-3}$ to $1\times 10^{-5}$ during epochs 1–2000, and keep $1\times 10^{-5}$ thereafter untill epoch 3000. We have total $200\text{k}$ samples and the batch size is $160\text{k}$ (full batch). All experiments are performed in JAX [50] (64–bit precision) using Equinox [59] modules and Optax optimizers; hyperparameters are tuned with Optuna [60]. Figure 10 reports the training and validation trajectories for the total loss together with the raw component losses from Eq. (31) (plotted before multiplying by their weights $\lambda_{\mathrm{\tiny{DynRes}}}{=}0.5$, $\lambda_{\tiny{J}}{=}1$, $\lambda_{\tiny{R}}{=}1$, $\lambda_{\tiny{Cons}}{=}0.1$, $\lambda_{\tiny{\dot{H}}}{=}0.1$). For monitoring, we also track a Hamiltonian loss (against the analytical $H$) that does not contribute to optimization.
Two components, the conservative–invariance term $\mathcal{L}_{\mathrm{\tiny{Cons}}}$ and the energy–rate consistency term $\mathcal{L}_{\tiny{\dot{H}}}$, start near zero, rise, and then descend alongside the others. This behaviour is expected as at initialization, the network can satisfy constraints involving $\nabla H_{\theta}$ with a nearly flat Hamiltonian (so $\nabla H_{\theta}\!\approx\!0$), yielding trivially small $\mathcal{L}_{\mathrm{\tiny{Cons}}}$ and $\mathcal{L}_{\tiny{\dot{H}}}$. As the dynamics and non-conservative terms ($\mathcal{L}_{\mathrm{\tiny{DynRes}}}$, $\mathcal{L}_{\tiny R}$) pull $H_{\theta}$ away from this degenerate solution, $\nabla H_{\theta}$ becomes nonzero; these losses temporarily increase and then decrease as all constraints are reconciled. Overall, the total loss decreases smoothly with a small train–validation gap, and we keep the checkpoint with the best validation loss and stop training once the validation loss fails to improve beyond a small sensitivity threshold for several epochs.
@@
Figure 11: Comparison of analytically derived quantities and those discovered by the pH–PINN. Panel (a) shows Hamiltonian trajectories; panel (b) shows the time derivative of the Hamiltonian. The close agreement indicates the model has captured the underlying energy structure from data.
Figure 11 compares the learned and analytical energies. Panel (a) plots the pH–PINN Hamiltonian $H_{\theta}(\mathbf{e})$ against the analytical $H$ from Eq. (24), while panel (b) compares the pH–PINN energy rate from Eq. (30), with the analytical $\dot{H}$ in Eq. (25). The close alignment across trajectories shows that the model has recovered both the energy landscape and the $\mathbf{J}/\mathbf{R}$ structure directly from data, without using the analytical energy expressions during training.
@@ 6 Summary and conclusions
(i) Global asymptotic synchronization via Lyapunov analysis. For two diffusively coupled neurons we derived the linearized error system and proved global asymptotic stability of the synchronization manifold $\mathcal{M}$ using a quadratic Lyapunov function. By bounding the error dynamics and enforcing positivity of the leading principal minors (Sylvester’s criterion), we obtained verifiable sufficient conditions guaranteeing $\dot{V}\leq 0$ and, by Barbalat’s lemma, $\mathbf{e}(t)\to\mathbf{0}$ as $t\to\infty$. The analysis makes explicit how electrical coupling and intrinsic parameters jointly suppress transverse perturbations.
@@
(ii) Synchronization energy via Helmholtz decomposition. We decomposed the error vector field into conservative and dissipative components and derived a closed-form synchronization Hamiltonian $H$ and its energy-rate identity $\dot{H}$ for the linearized error flow. This yields a quantitative measure of the energetic cost of synchrony and clarifies the respective roles of interconnection, damping, and the memristive offset in shaping energy exchange.
@@
(iii) Learning energetics with a port–Hamiltonian PINN. We proposed a port–Hamiltonian physics-informed neural network (pH-PINN) that embeds the conservative/non-conservative structure into training through physically motivated losses. The model learns $(H_{\theta},\mathbf{J}_{\phi},\mathbf{R}_{\psi})$ from data while enforcing skew-symmetry of $\mathbf{J}_{\phi}$, diagonality of $\mathbf{R}_{\psi}$, and the energy balance $\dot{H}_{\theta}=-\nabla H_{\theta}^{\top}\mathbf{R}_{\psi}\nabla H_{\theta}$. Numerically, the learned Hamiltonian and energy rate closely match their analytical counterparts.
@@
Direct simulations confirm complete synchronization and show monotone decay of the synchronization energy. Parameter sweeps across $(k,\rho,m,g_{e})$ reveal consistent trends between Lyapunov and Hamiltonian diagnostics: stronger electrical coupling accelerates synchronization, whereas excessive flux coupling can inject energy and delay convergence. Heatmaps of $\dot{V}$ and $\dot{H}$ agree across the parameter planes, supporting the equivalence of the two perspectives.
@@
Our proof and learned structure are derived for the linearized error dynamics in a two-neuron setting with a separable-gradient prior. Extending the analysis to (i) nonidentical neurons and parameter mismatch, (ii) networks with arbitrary topology, (iii) stochastic forcing and measurement noise, and (iv) partial observability and latent-state identification are natural next steps. Methodologically, lifting the separability prior via adaptive masking, handling the piecewise memristive nonlinearity beyond linearization, and incorporating uncertainty quantification into the pH-PINN would broaden applicability. Finally, deploying the framework on experimental data and leveraging it for structure-preserving control design (e.g., energy shaping and damping injection) are promising directions.
@@
In summary, the paper provides (a) verifiable synchronization conditions for a biophysically enriched Hindmarsh–Rose model, (b) a closed-form synchronization energy and its energy–rate law, and (c) a principled learning architecture that discovers the same energetic structure directly from data. Together, these results bridge rigorous dynamical-systems analysis with physics-informed machine learning for nonlinear neuronal synchronization and offer a template for energy-aware modeling in broader classes of complex systems.

### v3

@@ 4 Numerical simulations and discussion
To corroborate the stability results of Section 3.1 and the Hamiltonian-based analysis of Section 3.2, we performed numerical simulations of the synchronization-error dynamics in Eq. (6) simultaneously with of the coupled memristive HR neuron system in Eq. (2). The two neurons were initialized with slightly different values in all five state variables to induce synchronization transients. We then computed the time evolution of the error variables $(e_{x},e_{y},e_{z},e_{u},e_{\phi})$, together with the Lyapunov and synchronization Hamiltonian time derivatives $\dot{V}$ and $\dot{H}$, and analyzed their dependence on the parameters $m$, $k$, $\rho$, and $g_{e}$.
Figures 5(a2)-(e2) and (a1)-(e2) present the time series of the five synchronization–error components of the full nonlinear and linearized systems in Eqs. (5) and (6), respectively, each representing the difference between corresponding variables of the two coupled neurons. All error trajectories decay asymptotically to zero after a brief transient, confirming complete synchronization of the coupled system in Eq. (2). One notice that the fast variables $e_{x},e_{y},e_{u},e_{\phi}$ converge earlier than the slow adaptation variable $e_{z}$, reflecting the intrinsic separation of time scales in the neuron dynamics. These results corroborate the linearized transverse stability established in Section 3.1.
Figure 5: Time evolution of the five synchronization-error components $(e_{x},e_{y},e_{z},e_{u},e_{\phi})$ of the full nonlinear error system Eq. (5) in (a1)-(e1) and the linearized error system Eq. (6) in (a2)-(e2), all showing convergence to zero.
Figure 6 presents the temporal evolution of the synchronization Hamiltonian $H$ and its derivative $\dot{H}$ associated with the errors in Fig. 5(a2)-(e2) . The Hamiltonian decays to zero as $t\to\infty$ (overall relaxation toward synchrony) while $\dot{H}$ can take both positive and negative values during the transient and then stabilizes near zero as the neurons approach synchrony. The derivative $\dot{H}$ exhibits transient oscillations, taking both negative and positive values before stabilizing near zero; this reflects intermittent energy exchange due to memristive and flux-related terms in the dynamics. After transients, $\dot{H}\to 0$ and the system settles on the synchronization manifold, consistent with an overall dissipative relaxation toward synchrony.
Figure 6: Time evolution of the synchronization Hamiltonian function (a) and its time derivative (b), associated with the errors shown in Fig. 5(a2)-(e2) ; as the system synchronizes, both the synchronization Hamiltonian and its derivative approach zero.
@@
Figure 8: Heatmaps of the time derivatives of the Lyapunov function (a1)–(a3) and the corresponding synchronization Hamiltonian function (b1)–(b3) versus the coupling $g_{e}$ and the system parameters $k$, $m$, and $\rho$, respectively. Across all parameter pairs, $\dot{V}$ and $\dot{H}$ display consistent qualitative trends, demonstrating a robust correspondence between Lyapunov-based and Hamiltonian-based stability indicators.
In summary, the numerical results across the tested parameter ranges support the analytical pictures from Sections 3.1 and 3.2: (1) the linearized transverse dynamics are stable under the derived sufficient conditions, and (2) the Hamiltonian function decays to zero as $t\to\infty$ and the Lyapunov and Hamiltonian formulations yielding consistent measures of stability and energetics. These findings provide a basis for the pH-PINN framework introduced in Section 5, which exploits the equivalence between Lyapunov and Hamiltonian formalisms to learn the underlying Hamiltonian structure directly from data, bridging analytical theory with machine-learning discovery of Hamiltonian -conserving and dissipative dynamical systems.
@@ 5 A port-Hamiltonian physics-informed neural network (pH-PINN) approach
To study synchronization Hamiltonian of the coupled HR neurons in Eq. (2) from a data-driven perspective, we employ a pH–PINN. This framework is well-suited to our system, whose synchronization behavior admits a Hamiltonian structure consistent with the pH formulation of conservative and dissipative flows. Analytical Hamiltonian approaches can establish stability but often become infeasible for high-dimensional or partially known systems. In such settings, the governing physical laws, such as energy conservation and dissipation, remain valid even when the explicit equations are unknown. The pH–PINN incorporates these physical constraints in the learning process, allowing the synchronization Hamiltonian and its dynamics to be inferred directly from data. Here, the known analytical Hamiltonian serves as a benchmark to quantitatively validate the learned representation. In the sequel, we briefly outline the principles of PINNs, HNNs, and pHNNs, and describe how these are integrated to construct the proposed pH-PINN framework.
@@ 5.2 Hamiltonian neural networks (HNNs)
HNNs [36] employ a scalar-valued network $H_{\theta}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\!\to\!\mathbb{R}$ that maps the state vector $\mathbf{x}$ to a Hamiltonian $H_{\theta}(\mathbf{x})$. Automatic differentiation provides the gradient $\nabla H_{\theta}(\mathbf{x})$, representing partial derivatives of $H_{\theta}$ with respect to the state components. Rather than predicting the vector field directly, the dynamics are recovered through a (possibly state-dependent) skew-symmetric matrix $\mathbf{J}(\mathbf{x})\in\mathbb{R}^{n\times n}$:
@@ 5.3 Port-Hamiltonian neural networks (pHNN)
Here, $\mathbf{J}$ encodes power-conserving interconnections, and $\mathbf{R}\in\mathbb{R}^{n\times n}$ represents a non-conservative part of the dynamics. The Hamiltonian rate satisfies
@@ 5.4 Port–Hamiltonian physics-informed neural network (pH-PINN)
Each term enforces a distinct physical law: $\mathcal{L}_{\mathrm{\tiny DynRes}}$ ensures dynamic consistency; $\mathcal{L}_{\tiny J}$ and $\mathcal{L}_{\tiny R}$ align learned and analytical conservative/non-conservative flows; $\mathcal{L}_{\mathrm{\tiny Cons}}$ enforces Hamiltonian invariance along conservative trajectories; and $\mathcal{L}_{\tiny{\dot{H}}}$ enforces the Hamiltonian –rate identity (Eq. 46).
The synchronization Hamiltonian is parameterized by a multilayer perceptron (MLP) $f_{\theta}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\!\to\!\mathbb{R}$ and anchored at the synchronization manifold to remove gauge freedom and ensure a strict local minimum:
@@
This structure, inferred from $\mathbf{F}_{c}$ and $\mathbf{F}_{d}$ (Eq. (37)), reveals a sparse coupling pattern. We thus mask $\mathbf{J}_{\!\phi}$ and constrain $\mathbf{R}_{\!\psi}$ to be diagonal, consistent with standard pH learning [40]. Because the sparsity pattern depends on the Hamiltonian gradient $\nabla H$, we adopt a separable-gradient assumption, noting that many physical systems possess approximately separable Hamiltonian functions [36, 58, 38]. Each component $\partial H/\partial e_{i}$ thus depends primarily on its own coordinate (and possibly a small subset), consistent with the weak cross-partials observed in $\mathbf{F}_{c}$ and $\mathbf{F}_{d}$. Under this assumption, rows 2255 of $\mathbf{F}_{c}$ depend only on $e_{x}$, while $F_{c,x}$ depends on $e_{y},e_{u},$ and $e_{\phi}$ (but not $e_{z}$), motivating a masked, skew-symmetric $\mathbf{J}_{\!\phi}$ ($\mathbf{J}_{\!\phi}^{\!\top}=-\mathbf{J}_{\!\phi}$). Inspection of $\mathbf{F}_{d}$ further reveals primarily self-coupling (coordinate-wise damping) terms, motivating a diagonal parameterization of $\mathbf{R}_{\!\psi}$.
@@ 5.5 Training and Results
We train a pH–PINN by minimizing the total loss in Eq. (47). The ground-truth derivatives needed for the supervised learning are computed analytically by evaluating the model (Eq. (6)) right-hand sides at the states saved during ODE integration (no numerical/finite-difference differentiation is used on these states). Concretely, after integrating the coupled HR system in Eq. (2) together with the linearized error dynamics in Eq. (6) and storing the states on the solver time grid $\{t_{k}\}_{k=1}^{N}$, we compute $\dot{\mathbf{e}}^{\mathrm{true}}(t_{k})$ by direct substitution of the saved states into the right-hand side of Eq. (6) (and similarly for the decomposition $\mathbf{F}=\mathbf{F}_{c}+\mathbf{F}_{d}$ in Eq. (37)).
Each sample comprises the error state $\mathbf{e}\in\mathbb{R}^{5}$, its ground–truth derivative $\dot{\mathbf{e}}^{\mathrm{true}}$, the full neuron states of both units (used only to build a short context vector $\mathbf{c}=(x_{1},u_{1},u_{2},\phi_{1})^{\top}\in\mathbb{R}^{4}$), and the analytical Hamiltonian and Hamiltonian rate $(H,\dot{H})$ from Eqs. (40) and (41) for monitoring. The context vector $\mathbf{c}$ is required because the analytical conservative/dissipative fields in Eq. (37) are state-dependent through coefficients such as $N(x_{1},u_{1},\phi_{1})$ and the memristive switching terms $\alpha(u_{1},u_{2})$, $\beta(u_{1},u_{2})$; accordingly, the learned pH–PINN must be conditioned on $(x_{1},u_{1},u_{2},\phi_{1})$ to represent the same non-autonomous dynamics. Importantly, $\mathbf{c}$ contains only observed state variables needed to parameterize the right-hand side and does not include any supervision targets e.g., $\dot{\mathbf{e}}^{\mathrm{true}}$, $H$, or $\dot{H}$).
All quantities are kept in physical units and no normalization is used: applying a state normalization would require a consistent rescaling of $\mathbf{e}$, $\dot{\mathbf{e}}$, and the learned pH factors $(\mathbf{J},\mathbf{R},H)$ to preserve the scale of the induced prediction $\dot{\mathbf{e}}_{\mathrm{pred}}=(\mathbf{J}-\mathbf{R})\nabla H$ in the residual loss $\|\dot{\mathbf{e}}^{\mathrm{true}}-\dot{\mathbf{e}}_{\mathrm{pred}}\|^{2}$. Given that the simulated trajectories are bounded over the training window, we found training stable without additional scaling. All networks consume the concatenated $\mathbf{e},\mathbf{c}]$ as input. The overall architecture is shown in Fig. 9, and the implementation details are summarized in Table 1.
Figure 9: Schematic diagram of the proposed pH-PINN architecture.
@@
The dataset consists of N=200, [^0] N=200{,}000 time samples $\{(\mathbf{e}(t_{k}),\mathbf{c}(t_{k}),\dot{\mathbf{e}}^{\mathrm{true}}(t_{k}))\}_{k=1}^{N}$. We randomly permute the samples once and perform a fixed $80/20$ split into a training set and a held-out validation set (used only for evaluation and model selection). An epoch denotes one full pass through the training set. We use a batch size of 160,000160{,}000, which in our setup matches the training-set size after the split, so each epoch corresponds to a single full-batch AdamW update (more generally, the number of batches per epoch is $\lceil N_{\mathrm{train}}/\text{batch\_size}\rceil$). At the end of every epoch, we evaluate the total loss (and its components) on the entire validation set.
We train using the AdamW optimizer [59] with a linear learning-rate decay from $1\times 10^{-3}$ to $1\times 10^{-5}$ during epochs 1–2000, and keep $1\times 10^{-5}$ thereafter until epoch 3000. All experiments are performed in JAX [51] (64–bit precision) using Equinox [60] modules and Optax optimizers; hyperparameters are tuned with Optuna [61]. Figure 10 reports the training and validation trajectories for the total loss together with the raw component losses from Eq. (47) (plotted before multiplying by their weights $\lambda_{\mathrm{\tiny{DynRes}}}{=}0.5$, $\lambda_{\tiny{J}}{=}1$, $\lambda_{\tiny{R}}{=}1$, $\lambda_{\tiny{Cons}}{=}0.1$, $\lambda_{\tiny{\dot{H}}}{=}0.1$). For monitoring, we also track a Hamiltonian loss (against the analytical $H$) that does not contribute to optimization.
Two components, the conservative–invariance term $\mathcal{L}_{\mathrm{\tiny{Cons}}}$ and the Hamiltonian –rate consistency term $\mathcal{L}_{\tiny{\dot{H}}}$, start near zero, rise, and then descend alongside the others. This behavior is expected as at initialization, the network can satisfy constraints involving $\nabla H_{\theta}$ with a nearly flat Hamiltonian (so $\nabla H_{\theta}\!\approx\!0$), yielding trivially small $\mathcal{L}_{\mathrm{\tiny{Cons}}}$ and $\mathcal{L}_{\tiny{\dot{H}}}$. As the dynamics and non-conservative terms ($\mathcal{L}_{\mathrm{\tiny{DynRes}}}$, $\mathcal{L}_{\tiny R}$) pull $H_{\theta}$ away from this degenerate solution, $\nabla H_{\theta}$ becomes nonzero; these losses temporarily increase and then decrease as all constraints are reconciled. Overall, the total loss decreases smoothly with a small train–validation gap. Overall, the total loss decreases smoothly with a small train–validation gap. We retain the checkpoint achieving the lowest validation loss (evaluated at the end of each epoch on the full validation set) and report all results from this checkpoint. Training is run for a fixed budget of 3000 epochs (no early-stopping criterion is used).
@@
Figure 11: Comparison of analytically derived quantities and those discovered by the pH–PINN. Panel (a) shows Hamiltonian trajectories; panel (b) shows the time derivative of the Hamiltonian. The close agreement indicates the model has learned the underlying Hamiltonian structure from data.
Figure 11 compares the learned and analytical energies. Panel (a) plots the pH–PINN Hamiltonian $H_{\theta}(\mathbf{e})$ against the analytical $H$ from Eq. (40), while panel (b) compares the pH–PINN Hamiltonian rate from Eq. (46), with the analytical $\dot{H}$ in Eq. (41). The close alignment across trajectories shows that the model has recovered both the Hamiltonian landscape and the $\mathbf{J}/\mathbf{R}$ structure directly from data, without using the analytical expressions of the Hamiltonian during training.
@@ 6 Summary and conclusions
(i) Local synchronization via Lyapunov analysis. For two diffusively coupled neurons, we derived the linearized transverse error system and analyzed the stability of the synchronization manifold $\mathcal{M}$ associated with this linearized dynamics using a quadratic Lyapunov function. By bounding the time-varying coefficients along bounded reference trajectories and enforcing positivity of an explicit diagonal Lyapunov matrix, we obtained verifiable sufficient conditions for transverse stability. In dissipative memristive switching regimes, these conditions yield $\dot{V}\leq-\lambda_{\min}(\mathbf{M})\|\mathbf{e}\|^{2}$ and, via a Rayleigh bound and Barbalat’s lemma, asymptotic convergence $\|\mathbf{e}(t)\|\to 0$ (complete synchronization for the linearized error system). When the switching visits non-dissipative regimes, the same framework leads to a practical stability estimate of the form $\dot{V}\leq-2\alpha V+C$, implying uniform boundedness and an explicit ultimate bound on the synchronization error (practical synchronization). The analysis makes explicit how electrical coupling and intrinsic parameters jointly suppress transverse perturbations. A fully global nonlinear synchronization proof would require explicit bounds on the neglected nonlinear remainder terms, which we leave for future work.
@@
(ii) Synchronization Hamiltonian via Helmholtz decomposition. We decomposed the error vector field into conservative and dissipative components and derived a closed-form synchronization Hamiltonian $H$ and its rate identity $\dot{H}$ for the linearized error flow. This yields a quantitative measure of the energetic cost of synchrony and clarifies the respective roles of interconnection, damping, and the memristive offset in shaping energy exchange.
@@
(iii) Learning the synchronization Hamiltonian and its time rate with a port–Hamiltonian PINN. We proposed a port–Hamiltonian physics-informed neural network (pH-PINN) that embeds the conservative/non-conservative structure into training through physically motivated losses. The model learns $(H_{\theta},\mathbf{J}_{\phi},\mathbf{R}_{\psi})$ from data while enforcing skew-symmetry of $\mathbf{J}_{\phi}$, diagonality of $\mathbf{R}_{\psi}$, and the Hamiltonian balance $\dot{H}_{\theta}=-\nabla H_{\theta}^{\top}\mathbf{R}_{\psi}\nabla H_{\theta}$. Numerically, the learned Hamiltonian and its rate closely match their analytical counterparts.
@@
Direct simulations confirm complete synchronization and show decay of the synchronization Hamiltonian to zero as $t\to\infty$ . Parameter sweeps across $(k,\rho,m,g_{e})$ reveal consistent trends between Lyapunov and Hamiltonian diagnostics: stronger electrical coupling accelerates synchronization, whereas excessive flux coupling can inject energy and delay convergence. Heatmaps of $\dot{V}$ and $\dot{H}$ agree across the parameter planes, supporting the equivalence of the two perspectives.
@@
Our proof and learned structure are derived for the linearized error dynamics in a two-neuron setting with a separable-gradient prior. Extending the analysis to (a) nonidentical neurons and parameter mismatch, (b) larger networks with arbitrary topology, (c) stochastic forcing and measurement noise, and (d) partial observability and latent-state identification are natural next steps. Methodologically, lifting the separability prior via adaptive masking, handling the piecewise memristive nonlinearity beyond linearization, and incorporating uncertainty quantification into the pH-PINN would broaden applicability. Morever, deploying the framework on experimental data and leveraging it for structure-preserving control design (e.g., energy shaping and damping injection) are promising directions.
@@
In summary, the paper provides (I) verifiable conditions for synchronization in a biophysically enriched Hindmarsh–Rose model, (II) a closed-form synchronization Hamiltonian function together with an explicit rate law for its evolution, and (III) a principled learning architecture that recovers an equivalent Hamiltonian-based structure from data (up to scaling and additive constants). Collectively, these results connect rigorous dynamical systems analysis with physics-informed machine learning for nonlinear neuronal synchronization and provide a template for a port-Hamiltonian-aware modeling in broader classes of complex systems.